Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total and per unit

Sales (7,600 units) $258,400 $34.00

Variable expenses 144,400 19.00

Contribution margin 114,000 $15.00

Fixed expenses 55,300

Net operating income $58,700

Required: (Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 100 units?

2. What would be the revised net operating income per month if the sales volume decreases by 100 units?

3. What would be the revised net operating income per month if the sales volume is 6,600 units?
Answer #1

Net operating income = Total Contribution margin - Fixed cost

1. Net operating income = (15*7700) - 55,300

= 60,200

2. Net operating income = (15*7500) - 55,300

= 57,200

3. Net operating income = (15*6600) - 55,300

= 43,700